# A Near-optimal, Scalable and Corruption-tolerant Framework for Stochastic Bandits: From Single-Agent to Multi-Agent and Beyond

### v1

@@ 3 Stochastic Multi-Armed Bandits Robust to Adversarial Corruptions
1
2Initialization: Set the initial round $T_{0}=0$, $\Delta_{k}^{0}=1$, and $r_{k}^{0}=0$ for all $k\in[K]$.
3for epochs $m=1,2,\cdots$ do
4       Set $\zeta_{m}\leftarrow(m+4)2^{2(m+4)}\ln(K)$, and $\delta_{m}\leftarrow 1/(K\zeta_{m})$
5      Set $\lambda_{m}\leftarrow 2^{8}\ln{\left(4K/\delta_{m}\right)}$ and $\beta_{m}\leftarrow\delta_{m}/K$.
6      Set $n_{k}^{m}=\lambda_{m}(\Delta_{k}^{m-1})^{-2}$ for all arms $k\in[K]$.
7      Set $N_{m}\leftarrow\lceil K\lambda_{m}2^{2(m-1)}\rceil$ and $T_{m}\leftarrow T_{m-1}+N_{m}$.
8      Select the arm $k_{m}=\mathop{\arg\max}_{k\in[K]}r_{k}^{m-1}$.
9      Set $\widetilde{n}_{k}^{m}=\begin{cases}n_{k}^{m}&k\neq k_{m}\\ N_{m}-\sum_{k\neq k_{m}}n_{k}^{m}&k=k_{m}\end{cases}$.
10      for $t=T_{m-1}+1$ to $T_{m}$ do
11             Choose arm $I_{t}\sim p_{m}$ where $p_{m}(k)=\widetilde{n}_{k}^{m}/N_{m}$.
12            Observe the corrupted reward $\widetilde{r}_{t,I_{t}}$ and update the total reward $S_{I_{t}}^{m}=S_{I_{t}}^{m}+\widetilde{r}_{t,I_{t}}$.
13       end for
14
15      Set $r_{k}^{m}\leftarrow\min\{S_{k}^{m}/\widetilde{n}_{k}^{m},1\}$
[…]
17      Set $\Delta_{k}^{m}\leftarrow\max\{2^{-m},r_{*}^{m}-r_{k}^{m}\}$.
18 end for
@@ 15      Set $r_{k}^{m}\leftarrow\min\{S_{k}^{m}/\widetilde{n}_{k}^{m},1\}$
We introduce our BARBAT framework for stochastic bandits with adversarial corruptions in Algorithm 1. In each epoch $m$, BARBAT sets $\delta_{m}=1/(K\zeta_{m})$ where $\zeta_{m}$ is chosen to make $\delta_{m}\leq 1/N_{m}$. From $\delta_{m}$, BARBAT set parameters $\lambda_{m}$ and $\beta_{m}$ based on $\delta_{m}$. It then assigns $n_{k}^{m}=\lambda_{m}\bigl{(}\Delta_{k}^{m-1}\bigr{)}^{-2}$, where $\Delta_{k}^{0}=1$. Unlike BARBAR, which defines $N_{m}=\sum_{k\in[K]}\lambda_{m}n_{k}^{m}$, BARBAT employs $N_{m}=\bigl{\lceil}K\lambda_{m}2^{2(m-1)}\bigr{\rceil}\geq\sum_{k\in[K]}n_{k}^{m}.$ It selects the arm with the highest empirical reward in the previous epoch $k_{m}=\arg\max_{k\in[K]}r_{k}^{m-1}$ (if multiple arms has the same mean rewards, we select the arm with the smallest index) and sets
$\widetilde{n}_{k}^{m}=\begin{cases}n_{k}^{m},&k\neq k_{m},\\ N_{m}-\sum_{k\neq k_{m}}n_{k}^{m},&k=k_{m}.\end{cases}$
During epoch $m$, the agent pulls arm $I_{t}$ with probability $p_{m}(I_{t})=\widetilde{n}_{I_{t}}^{m}/N_{m}$, observes the corrupted reward $\widetilde{r}_{t,I_{t}}$, and accumulates $S_{I_{t}}^{m}$. At the end of each epoch $m$, BARBAT sets $r_{k}^{m}=\min\{S_{k}^{m}/\widetilde{n}_{k}^{m},1\}$, ensuring $r_{k}^{m}\leq 1$ and addressing the BARBAR issue noted by Lu et al. (2021). The algorithm estimates $r_{*}^{m}$ and updates the estimated suboptimality gap $\Delta_{k}^{m}$.
The key difference between BARBAT and BARBAR is that epoch length of BARBAR heavily depends on the empirical rewards, so that adversary may manipulate the epoch length and potentially increase the chance of pulling suboptimal arms (see Appendix B of Gupta et al. (2019)). On the other hand, our BARBAT algorithm always guarantees that $N_{m+1}/N_{m}\in[4,14/3]$ and makes the algorithm more robust. By our choice of $N_{m}$ and $\delta_{m}$, we can reduce the regret bound from $O(KC)$ to $O(C)$ with the cost of an extra term $O\bigl{(}K\log^{2}(1/\Delta)/\Delta\bigr{)}$, which is independent of $T$. Moreover, BARBAT is an anytime algorithm, i.e., it does not require prior knowledge of $T$. BARBAT achieves the following regret bound, with the proof provided in Appendix A:
Theorem 1
@@
${\mathbb{E}}[R(T)]=O\left(C+\sum_{\Delta_{k}>0}\frac{\log(T)\log(KT)}{\Delta_{k}}+\frac{K\log\left(1/\Delta\right)\log\left(K/\Delta\right)}{\Delta}\right).$
Remark 1
BARBAT removes the multiplicative factor $K$ in the corruption-dependent term, addressing the open problem of Gupta et al. (2019). Our method also outperform the concurrent work (Ghaffari et al., 2024), which incur $O\bigl{(}C+K\log^{2}(T)/\Delta\bigr{)}$ regret.
@@ 4 Extensions
In this section, we extend BARBAT to various corrupted bandit settings discussed earlier and introduce necessary modifications to accommodate the relevant environment configurations.
@@ 4.1 Strongly Observable Graph Bandits
In strongly observable graph bandits, the feedback structure is defined by a directed graph $G=(K,E)$, where pulling arm $k_{i}$ reveals rewards for all arms $k_{j}$ such that $(k_{i},k_{j})\in E$. The independence number $\alpha$ is the size of the largest independent set in $G$. For the strongly observable graphs, each arm has either a self-loop or incoming edges from all other arms, revealing multiple rewards in a single pull (Alon et al., 2017).
Most previous works (Ito et al., 2022; Dann et al., 2023; Eldowa et al., 2024) rely on the knowledge of independence number $\alpha$, which is NP-hard to compute. Ito et al. (2024) addresses this in undirected graph bandits with self-loops but assumes a unique optimal arm.
To overcome these issues, we extend BARBAT to SOG-BARBAT (Algorithm 2). In epoch $m$, we observe $\lambda_{m}(\Delta_{k}^{m-1})^{-2}$ times for arm $k$ in expected. SOG-BARBAT aims to pull the fewest arms while ensuring sufficient observations. It defines expected pulling times array $Z^{m}[K]$ and expected observable times array $H^{m}[K]$. For ease of analysis, we assume that each arm is equipped with a self-loop. From a theoretical standpoint, this assumption does not alter the resulting regret. We design the OODS algorithm (Algorithm 3), which ensures that the out-dominating set $D_{s}^{m}$ satisfies $|D_{s}^{m}|\leq\alpha(1+2\ln(K/\alpha))$. For acyclic graphs, we always have $D_{s}^{m}\leq\alpha$. We compute ${\mathcal{H}}_{s}^{m}=\min_{k\in[K]}$ for each $k_{i}\in D_{s}^{m}$, set $Z_{k_{i}}^{m}=Z_{k_{i}}^{m}+{\mathcal{H}}_{s}^{m}$, and update $H^{m}$. Repeating these steps ensures that all arms are observed sufficiently during epoch $m$.
1
Input: A Strongly Observable Directed Graph $G$
2
3Initialization: Set the initial round $T_{0}=0$, $\Delta_{k}^{0}=1$, and $r_{k}^{0}=0$ for all $k\in[K]$.
4for epochs $m=1,2,...$ do
5       Set $\zeta_{m}\leftarrow(m+4)2^{2(m+4)}\ln(K)$, and $\delta_{m}\leftarrow 1/(K\zeta_{m})$
6      Set $\lambda_{m}\leftarrow 2^{8}\ln{\left(4K/\delta_{m}\right)}$ and $\beta_{m}\leftarrow\delta_{m}/K$.
7      Set $n_{k}^{m}=\lambda_{m}(\Delta_{k}^{m-1})^{-2}$ for all arms $k\in[K]$.
8      Set $N_{m}\leftarrow\lceil K\lambda_{m}2^{2(m-1)}\rceil$ and $T_{m}\leftarrow T_{m-1}+N_{m}$.
9      Select the arm $k_{m}=\mathop{\arg\max}_{k\in[K]}r_{k}^{m-1}$.
10      Build two arrays $Z_{k}^{m}\leftarrow 0$ and $H_{k}^{m}\leftarrow 0$ for all arms $k\in[K]$.
11      Set $s=1$, Copy graph $G$ to obtain a new graph $G_{s}^{m}$.
12      while $H_{k}^{m}\geq n_{k}^{m}$ holds for all arms $k\in[K]$ do
13
14            Run Algorithm 3 with parameter $G_{s}^{m}$ to obtain an out-dominating set $D_{s}^{m}$.
15            Set ${\mathcal{H}}_{s}^{m}=\min_{k\in[K]}(n_{k}^{m}-H_{k}^{m})$.
16            For each arm $k_{i}\in D_{s}^{m}$, set $Z_{k_{i}}^{m}=Z_{k_{i}}^{m}+{\mathcal{H}}_{s}^{m}$, and $H_{k_{j}}^{m}=H_{k_{j}}^{m}+{\mathcal{H}}_{s}^{m}$ for all arms $(k_{i},k_{j})\in E$ and self $k_{i}$, set $s\leftarrow s+1$.
17            Remove all arms which satisfies $H_{k}^{m}\geq n_{k}^{m}$ to obtain a new graph $G_{s}^{m}$.
18       end while
19
20      Set $\widetilde{n}_{k}^{m}=\begin{cases}Z_{k}^{m}&k\neq k_{m}\\ N_{m}-\sum_{k\neq k_{m}}Z_{k}^{m}&k=k_{m}\end{cases}$, and $\hat{n}_{k_{j}}^{m}=\sum_{(k_{i},k_{j})\in E}\widetilde{n}_{k_{i}}^{m}$.
21      for $t=T_{m-1}+1$ to $T_{m}$ do
22             Choose arm $I_{t}\sim p_{m}$ where $p_{m}(k)=\widetilde{n}_{k}^{m}/N_{m}$.
23            Observe the corrupted reward $\widetilde{r}_{t,I_{t}}$ and update the total reward $S_{I_{t}}^{m}=S_{I_{t}}^{m}+\widetilde{r}_{t,I_{t}}$.
24       end for
25
26      Set $r_{k}^{m}\leftarrow\min\{S_{k}^{m}/\hat{n}_{k}^{m},1\}$. Set $r_{*}^{m}\leftarrow\max_{k\in[K]}\left\{r_{k}^{m}-\sqrt{\frac{4\ln(4/\beta_{m})}{\hat{n}_{k}^{m}}}\right\}$
27      Set $\Delta_{k}^{m}\leftarrow\max\{2^{-m},r_{*}^{m}-r_{k}^{m}\}$.
28 end for
@@ 5       Set $\zeta_{m}\leftarrow(m+4)2^{2(m+4)}\ln(K)$, and $\delta_{m}\leftarrow 1/(K\zeta_{m})$
Input: A Strongly Observable Directed Graph $G$
Output: An Out-Dominating Set $D$
1 while the graph $G$ is empty do
2       if there exist cycles in graph $G$ then
3             Select the vertex with the largest out-degree in the graph $G$ and add it to the set $D$.
4            Remove the vertex and its out-degree neighbors from graph $G$.
5      else
6             Find all no-root vertices and add them to the set $D$.
7            Remove all no-root vertices and their out-degree neighbors from graph $G$.
8       end if
9
10 end while
Algorithm 3 OODS: Obtain an Out-Dominating Set
The regret bound for SOG-BARBAT as follows, the proof is given in Appendix B.
Theorem 2
For any strongly observable directed graph $G=(K,E)$ with independence number $\alpha$, the expected regret of SOG-BARBAT satisfies
${\mathbb{E}}\left[R(T)\right]=O\left(C+\sum_{k\in{\mathcal{I}}^{*}}\frac{\log(T)\log(KT)}{\Delta_{k}}+\frac{K\log\left(1/\Delta\right)\log\left(K/\Delta\right)}{\Delta}\right).$
For general strongly observable graphs, ${\mathcal{I}}^{*}$ is the set of at most $\lfloor\alpha(1+2\ln(K/\alpha))\rfloor$ arms with the smallest gaps. Specially, for acyclic graphs (include undirected graphs), ${\mathcal{I}}^{*}$ is the set of at most $\alpha$ arms with the smallest gaps.
Remark 2
The regret upper bound of SOG-BARBAT depends on $|{\mathcal{I}}^{*}|=O(\alpha\ln(K/\alpha+1))$, which matches the lower bound established by Chen et al. (2023). When $G$ is an acyclic graph, we can get $|{\mathcal{I}}^{*}|\leq\alpha$, indicating that SOG-BARBAT performs better under this condition.
Compared to recent works (Ito et al., 2022; Dann et al., 2023), whose regret bounds depend on the smallest suboptimality gap $1/\Delta$, our bound reveals the relationship between the regret and the suboptimality gap of all arms. Notably, these works require prior knowledge of the independence number $\alpha$ which is NP-hard to compute, whereas SOG-BARBAT does not need such prior knowledge.

### v2

@@ 3 Stochastic Multi-Armed Bandits Robust to Adversarial Corruptions
Main idea
Notice that in the BARBAR (gupta2019better) algorithm, the arm $k$ is approximately drawn $n_{k}^{m}=O(1/(\Delta_{k}^{m-1})^{2})$ times in expectation during epoch $m$, where $\Delta_{k}^{m-1}$ is the estimated gap computed in the previous epoch. Thus, each epoch has a length of approximately $N_{m}\approx O(\sum_{k}1/(\Delta_{k}^{m-1})^{2})$. Consequently, the adversary can significantly extend the length of an epoch by utilizing all of the corruption budget in the previous epoch, thereby implicitly increasing the number of pulls for suboptimal arms. To address this challenge, our BARBAT algorithm chooses a data-independent epoch length $N_{m}\approx O(K2^{2(m-1)})$, which cannot be affected by the adversary. We still allocate $n_{k}^{m}=O(1/(\Delta_{k}^{m-1})^{2})$ pulls to each arm $k$, but assign all remaining pulls to the estimated optimal arm $k_{m}$, resulting in pulling arm $k_{m}$ for about $N_{m}-\sum_{k\neq k_{m}}n_{k}^{m}$ rounds. On the other hand, pulling the estimated optimal arm $k_{m}$ too many times may lead to additional regret, as arm $k_{m}$ could be suboptimal. To mitigate this additional regret, we adopt epoch-varying failure probabilities $\delta_{m}\approx O(1/(mK2^{2m}))$ for epoch $m$ rather than using a global failure probability $\delta$. The choice of $\delta_{m}$ also eliminates the need to know the time horizon $T$.
@@
1:Initialization: Set the initial round $T_{0}\leftarrow 0$, $\Delta_{k}^{0}\leftarrow 1$, and $r_{k}^{0}\leftarrow 0$ for all $k\in[K]$.
2:for epochs $m=1,2,\cdots$ do
3:  Set $\zeta_{m}\leftarrow(m{+}4)2^{2(m+4)}\ln(K)$ and $\delta_{m}\leftarrow 1/(K\zeta_{m})$.
4:  Set $\lambda_{m}\leftarrow 2^{8}\ln{\left(4K/\delta_{m}\right)}$ and $\beta_{m}\leftarrow\delta_{m}/K$.
5:  Set $n_{k}^{m}\leftarrow\lambda_{m}(\Delta_{k}^{m-1})^{-2}$ for all $k\in[K]$.
6:  Set $N_{m}\leftarrow\lceil K\lambda_{m}2^{2(m-1)}\rceil$ and $T_{m}\leftarrow T_{m-1}+N_{m}$.
7:  Set $k_{m}\leftarrow\mathop{\arg\max}_{k\in[K]}r_{k}^{m-1}$ and compute $\widetilde{n}_{k}^{m}$ according to (1).
8:  for $t=T_{m-1}+1$ to $T_{m}$ do
9:   Choose arm $I_{t}\sim p_{m}$ where $p_{m}(k)=\widetilde{n}_{k}^{m}/N_{m}$.
10:   Observe the corrupted reward $\widetilde{r}_{t,I_{t}}$ and update the total reward $S_{I_{t}}^{m}\leftarrow S_{I_{t}}^{m}+\widetilde{r}_{t,I_{t}}$.
11:  end for
12:  Set $r_{k}^{m}\leftarrow\min\{S_{k}^{m}/\widetilde{n}_{k}^{m},1\}$ and $r_{*}^{m}\leftarrow\max_{k\in[K]}\Big\{r_{k}^{m}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}\Big\}$.
13:  Set $\Delta_{k}^{m}\leftarrow\max\{2^{-m},r_{*}^{m}-r_{k}^{m}\}$.
14:end for
Algorithm
We introduce our BARBAT framework for stochastic bandits with adversarial corruptions in Algorithm 1. In each epoch $m$, BARBAT sets the failure probability as $\delta_{m}=1/(K\zeta_{m})$, where $\zeta_{m}$ is chosen to make $\delta_{m}\leq 1/N_{m}$. The algorithm sets the epoch length to be $N_{m}=\big\lceil K\lambda_{m}2^{2(m-1)}\big\rceil\geq\sum_{k\in[K]}n_{k}^{m}$. We denote $k_{m}$ as the arm with the maximum empirical reward in the previous epoch; in case of ties, the arm with the smallest index is chosen. Then the number of pulls for each arm is set to
$\widetilde{n}_{k}^{m}=\begin{cases}n_{k}^{m},&k\neq k_{m},\\ N_{m}-\sum_{k\neq k_{m}}n_{k}^{m},&k=k_{m}.\end{cases}$ (1)
During epoch $m$, the agent pulls arm $I_{t}$ with probability $p_{m}(I_{t})=\widetilde{n}_{I_{t}}^{m}/N_{m}$, observes the corrupted reward $\widetilde{r}_{t,I_{t}}$, and accumulates the total reward $S_{I_{t}}^{m}$. At the end of each epoch $m$, BARBAT updates the estimated reward $r_{k}^{m}=\min\{S_{k}^{m}/\widetilde{n}_{k}^{m},1\}$ and computes the estimated suboptimality gap $\Delta_{k}^{m}$.
An example showing how BARBAT avoids paying an $O(KC)$ term
Gupta et al. gupta2019better provide an example where BARBAR suffers regret $\Omega(KC)$, in which all corruptions happen in an epoch $c$ with corruption level $C=N_{c}=\lambda(2^{2(c-1)}+O(K))$. Then the algorithm will lose all information from the previous epochs so that the length of the next epoch becomes $N_{c+1}=K\lambda 2^{2c}=O(KC)$ and each arm will be pulled uniformly. In this scenario, BARBAR will pull each arm approximately $O(C)$ times and generate O((K1)C)O((K-1)C) regret. On the other hand, our BARBAT algorithm fixes the length of the $m$-th epoch to be $N_{m}=K\lambda_{m}2^{2(m-1)}$. If the adversary intends to corrupt an entire epoch $c$, the required corruption level is $C=N_{c}=K\lambda 2^{2(c-1)}$, which is $K$ times larger than needed to attack BARBAR. Thus our BARBAT algorithm will only suffer $O(C)$ regret from adversarial corruptions.
Regret Bound
BARBAT achieves the following regret bound, with proofs deferred to Appendix B.1.
Theorem 1.
@@
$R(T)=O\bigg(C+\sum_{\Delta_{k}>0}\frac{\log(T)\log(KT)}{\Delta_{k}}+\frac{K\log(1/\Delta)\log(K/\Delta)}{\Delta}\bigg).$
Remark 1.
The regret bound of BARBAT contains an extra term $K\log^{2}(1/\Delta)/\Delta$. Notice that this term is independent of $T$, making it generally much smaller than the major term $\sum_{\Delta_{k}>0}\frac{\log^{2}(T)}{\Delta_{k}}$.
Remark 2.
Notice that a concurrent work ghaffari2025multi also employs static epoch lengths and removes the factor $K$ from the regret bound. However, they adopt a global $\delta$ across all epochs, resulting in a regret of $O(C+K\log^{2}(T)/\Delta)$, which is worse than our regret of $O\left(C+\sum_{\Delta_{k}>0}\frac{\log^{2}(T)}{\Delta_{k}}\right)$.
@@ 4 Extensions
In this section, we extend BARBAT to various corrupted bandit settings and introduce the necessary modifications to accommodate the specific configurations of these environments.
4.1 Cooperative Multi-Agent Multi-Armed Bandits
We extend BARBAT to the multi-agent setting and propose the MA-BARBAT (Multi-Agent BARBAT) algorithm in Algorithm 2. In each epoch $m$, all agents pull arms according to the same probability distribution and broadcast their total rewards $\{S_{v,k}^{m}\}_{k=1}^{K}$ at the end of the epoch. Each agent then updates the estimated suboptimality gaps $\Delta_{k}^{m}$ using the received messages and its own observations. We present the regret bound and communication cost of our MA-BARBAT algorithm in the following theorem, with the proof given in Appendix B.2.
Algorithm 2 MA-BARBAT
1:Initialization: Set the initial round $T_{0}\leftarrow 0$, $\Delta_{k}^{0}\leftarrow 1$, and $r_{k}^{0}\leftarrow 0$ for all $k\in[K]$.
2:for epoch $m=1,2,\ldots$ do
3:  Set $\zeta_{m}\leftarrow(m+4)2^{2(m+4)}\ln(VK)$ and $\delta_{m}\leftarrow 1/(VK\zeta_{m})$.
4:  Set $\lambda_{m}\leftarrow 2^{8}\ln\left(4K/\delta_{m}\right)/V$ and $\beta_{m}\leftarrow\delta_{m}/(VK)$.
5:  Set $n_{k}^{m}\leftarrow\lambda_{m}(\Delta_{k}^{m-1})^{-2}$ for all $k\in[K]$.
6:  Set $N_{m}\leftarrow\lceil K\lambda_{m}2^{2(m-1)}\rceil$ and $T_{m}\leftarrow T_{m-1}+N_{m}$.
7:  Set $k_{m}\leftarrow\mathop{\arg\max}_{k\in[K]}r_{k}^{m-1}$ and compute $\widetilde{n}_{k}^{m}$ according to (1).
8:  for $t=T_{m-1}+1$ to $T_{m}$ do
9:   Each agent chooses arm $I_{t}\sim p_{m}$, where $p_{m}(k)=\widetilde{n}_{k}^{m}/N_{m}$.
10:   Each agent observes the corrupted reward $\widetilde{r}_{t,I_{t}}$ and compute $S_{I_{t}}^{m}\leftarrow S_{I_{t}}^{m}+\widetilde{r}_{t,I_{t}}.$
11:  end for
12:  Each agent $v$ broadcasts the messages $\{S_{v,k}^{m}\}_{k\in[K]}$ to other agents.
13:  Set $r_{k}^{m}\leftarrow\min\{\sum_{v\in[V]}S_{v,k}^{m}/(V\widetilde{n}_{k}^{m}),1\}$ and $r_{*}^{m}{\leftarrow}\max_{k\in[K]}\Big\{r_{k}^{m}-\sqrt{\frac{4\ln(4/\beta_{m})}{V\widetilde{n}_{k}^{m}}}\Big\}$.
14:  Set $\Delta_{k}^{m}\leftarrow\max\{2^{-m},r_{*}^{m}-r_{k}^{m}\}$.
15:end for
Theorem 2.
The MA-BARBAT algorithm requires only a communication cost of $V\log(VT)$ over $T$ rounds. The individual regret of each agent $v$ satisfies:
$R_{v}(T)=O\left(\frac{C}{V}+\sum_{\Delta_{k}>0}\frac{\log(VT)\log(KVT)}{V\Delta_{k}}+\frac{K\log\left(1/\Delta\right)\log\left(KV/\Delta\right)}{V\Delta}\right).$
Remark 3.
The regret bound demonstrates that collaboration reduces each agent’s individual regret by a factor of $V$. As summarized in Table 1, our bound is tighter than those of previous works on CMA2B with adversarial corruptions.
Remark 4.
Notice that some recent works wang2023achieving on CMA2B achieve a communication cost of only $o(\log(T))$. However, these methods rely on the assumption of a unique best arm. In the presence of multiple best arms, it remains unclear whether a communication cost of $o(\log(T))$ is sufficient.
4.2 Batched Bandits
We extend BARBAT to the batched bandit setting and propose the BB-BARBAT (Batched Bandits-BARBAT) algorithm in Algorithm 3. The BB-BARBAT algorithm sets the epoch lengths $N_{m}$ to be approximately $O(T^{\frac{m}{L+1}})$ ensuring that the number of epochs matches the number of batches. In this way, each epoch serves as a batch, making the BB-BARBAT algorithm almost identical to the BARBAT algorithm.
Algorithm 3 $B$-BARBAT
1:Initialization: Set the initial round $T_{0}\leftarrow 0$, $\Delta_{k}^{0}\leftarrow 1$, and $r_{k}^{0}\leftarrow 0$ for all $k\in[K]$.
2:for epochs $m=1,2,\cdots$ do
3:  Set $a\leftarrow T^{\frac{1}{2(L+1)}}$, $\zeta_{m}\leftarrow(m+4)a^{2(m+4)}\ln(aK)$ and $\delta_{m}\leftarrow 1/(K\zeta_{m})$.
4:  Set $\lambda_{m}\leftarrow a^{8}\ln{\left(4K/\delta_{m}\right)}$ and $\beta_{m}\leftarrow\delta_{m}/K$.
5:  Set $n_{k}^{m}\leftarrow\lambda_{m}(\Delta_{k}^{m-1})^{-2}$ for all $k\in[K]$.
6:  Set $N_{m}\leftarrow\lceil K\lambda_{m}a^{2(m-1)}\rceil$ and $T_{m}\leftarrow T_{m-1}+N_{m}$.
7:  Set $k_{m}\leftarrow\mathop{\arg\max}_{k\in[K]}r_{k}^{m-1}$ and compute $\widetilde{n}_{k}^{m}$ according to (1).
8:  for $t=T_{m-1}+1$ to $T_{m}$ do
9:   Choose arm $I_{t}\sim p_{m}$ where $p_{m}(k)=\widetilde{n}_{k}^{m}/N_{m}$.
10:  end for
11:  Observe the all corrupted rewards $\widetilde{r}_{I_{t}}$ in this batch and compute $S_{I_{t}}^{m}\leftarrow S_{I_{t}}^{m}+\widetilde{r}_{I_{t}}$.
12:  Set $r_{k}^{m}\leftarrow\min\{S_{k}^{m}/\widetilde{n}_{k}^{m},1\}$ and $r_{*}^{m}{\leftarrow}\max_{k\in[K]}\left\{r_{k}^{m}-\sqrt{\frac{4\ln(4/\beta_{m})}{\widetilde{n}_{k}^{m}}}\right\}$.
13:  Set $\Delta_{k}^{m}\leftarrow\max\{a^{-m},r_{*}^{m}-r_{k}^{m}\}$.
14:end for
We present the regret upper bound of BB-BARBAT and the lower bound for corrupted batched bandits in Theorems 3 and 4, respectively. The proofs are deferred to Appendix B.3.
Theorem 3.
For corrupted batched bandits with a static grid, the regret of BB-BARBAT satisfies
$R(T)=O\bigg(CT^{\frac{1}{L+3}}+T^{\frac{4}{L+3}}\bigg(\sum_{\Delta_{k}>0}\frac{L\log(KT)}{\Delta_{k}}+\frac{K\log(T)\log(1/\Delta)\log(K/\Delta)}{L\Delta}\bigg)\bigg).$
Theorem 4.
For any algorithm, there exists an instance of batched bandits with a static grid such that
$R(T)\geq\Omega\left(T^{\frac{1}{L}}\left(K+C^{1-\frac{1}{L}}\right)\right).$
Remark 5.
There is still a gap between our upper bound and the lower bound. We believe that our lower bound is nearly tight. Achieving optimal regret upper bounds for corrupted batched bandits remains an interesting open question.
@@ 4.3 Strongly Observable Graph Bandits
We extend BARBAT to strongly observable graph bandits and propose the SOG-BARBAT (Strongly-Observable Graph bandits-BARBAT) algorithm, presented in Algorithm 4. Our method adopts $Z^{m}$ and $H^{m}$ to denote the expected pulls and expected observations in epoch $m$, respectively. $G^{m}$ denotes the graph used in the $m$-th epoch. SOG-BARBAT adopts the proposed Algorithm 6 (in Appendix A) to compute the out-domination set $D^{m}$ of $G^{m}$, whose size is at most $\alpha(1+2\ln(K/\alpha))$. Moreover, if $G^{m}$ is acyclic, we can guarantee that $|D^{m}|\leq\alpha$. Then SOG-BARBAT computes $\overline{H}^{m}$, which is the minimal number of additional pulls per out-dominating arm that ensures some vertex meets its observation requirement. After that, SOG-BARBAT updates $Z^{m}$ and $H^{m}$, and removes all nodes that have now been sufficiently observed from $G^{m}$. Repeating these steps ensures that by the end of epoch $m$, all arms have been observed sufficiently.
@@
1:Input: A Strongly Observable Directed Graph $G$.
2:Initialization: Set the initial round $T_{0}\leftarrow 0$, $\Delta_{k}^{0}\leftarrow 1$, and $r_{k}^{0}\leftarrow 0$ for all $k\in[K]$.
3:for epochs $m=1,2,$... do
4:  Set $\zeta_{m}\leftarrow(m+4)2^{2(m+4)}\ln(K)$ and $\delta_{m}\leftarrow 1/(K\zeta_{m})$.
5:  Set $\lambda_{m}\leftarrow 2^{8}\ln{\left(4K/\delta_{m}\right)}$ and $\beta_{m}\leftarrow\delta_{m}/K$.
6:  Set $n_{k}^{m}\leftarrow\lambda_{m}(\Delta_{k}^{m-1})^{-2}$ for all $k\in[K]$.
7:  Set $N_{m}\leftarrow\lceil K\lambda_{m}2^{2(m-1)}\rceil$ and $T_{m}\leftarrow T_{m-1}+N_{m}$.
8:  Set $k_{m}\leftarrow\mathop{\arg\max}_{k\in[K]}r_{k}^{m-1}$.
9:  Set $Z_{k}^{m}\leftarrow 0$ and $H_{k}^{m}\leftarrow 0$ for all arms $k\in[K]$, $G^{m}\leftarrow G$.
10:  while $H_{k}^{m}\geq n_{k}^{m}$ holds for all arms $k\in[K]$ do
11:   Compute an out-domination set $D^{m}$ of $G^{m}$ by Algorithm 6.
12:   Compute $\overline{H}^{m}\leftarrow\min_{k\in[K]}(n_{k}^{m}-H_{k}^{m})$.
13:   for each arm $k_{i}\in D^{m}$ do
14:    Update $Z^{m}_{k_{i}}\leftarrow Z^{m}_{k_{i}}+\overline{H}_{s}^{m}$ and $H^{m}_{k_{i}}\leftarrow H^{m}_{k_{i}}+\overline{H}^{m}$
15:    Update $H^{m}_{k_{j}}\leftarrow H^{m}_{k_{j}}+\overline{H}^{m}$ for all arms $(k_{i},k_{j})\in E$ with $k_{j}\neq k_{i}$
16:   end for
17:   Remove all arms which satisfy $H_{k}^{m}\geq n_{k}^{m}$ from $G^{m}$.
18:  end while
19:  Set $\widetilde{n}_{k}^{m}=\begin{cases}Z_{k}^{m}&k\neq k_{m}\\ N_{m}-\sum_{k\neq k_{m}}Z_{k}^{m}&k=k_{m}\end{cases}$, and $\hat{n}_{k_{j}}^{m}\leftarrow\sum_{(k_{i},k_{j})\in E}\widetilde{n}_{k_{i}}^{m}$.
20:  for $t=T_{m-1}+1$ to $T_{m}$ do
21:   Choose arm $I_{t}\sim p_{m}$ where $p_{m}(k)=\widetilde{n}_{k}^{m}/N_{m}$.
22:   Observe the corrupted reward $\widetilde{r}_{t,I_{t}}$ and update the total reward $S_{I_{t}}^{m}\leftarrow S_{I_{t}}^{m}+\widetilde{r}_{t,I_{t}}$.
23:  end for
24:  Set $r_{k}^{m}\leftarrow\min\{S_{k}^{m}/\hat{n}_{k}^{m},1\}$ and $r_{*}^{m}\leftarrow\max_{k\in[K]}\left\{r_{k}^{m}-\sqrt{\frac{4\ln(4/\beta_{m})}{\hat{n}_{k}^{m}}}\right\}$
25:  Set $\Delta_{k}^{m}\leftarrow\max\{2^{-m},r_{*}^{m}-r_{k}^{m}\}$.
26:end for
The regret bound for SOG-BARBAT is presented as follows, with the proof given in Appendix B.4.
Theorem 5.
For any strongly observable directed graph $G=([K],E)$ with independence number $\alpha$, the expected regret of SOG-BARBAT satisfies
${\mathbb{E}}\left[R(T)\right]=O\left(C+\sum_{k\in{\mathcal{I}}^{*}}\frac{\log(T)\log(KT)}{\Delta_{k}}+\frac{K\log\left(1/\Delta\right)\log\left(K/\Delta\right)}{\Delta}\right),$
where $\mathcal{I}$$^{*}$ is the set of at most $\lceil\alpha(1+2\ln(K/\alpha))\rceil$ arms with the smallest gaps. Especially, for directed acyclic graphs and undirected graphs, $\mathcal{I}$$^{*}$ is the set of at most $\alpha$ arms with the smallest gaps.
Remark 6.
The regret upper bound of SOG-BARBAT depends on $|{\mathcal{I}}^{*}|=O(\alpha\ln(K/\alpha+1))$, which matches the lower bound established by chen2024interpolating. When $G$ is an acyclic graph or an undirected graph, we have $|{\mathcal{I}}^{*}|\leq\alpha$, indicating that SOG-BARBAT performs better under this condition.
Remark 7.
Compared to FTRL-based methods (ito2022nearly; dann2023blackbox), whose regret bounds depend on the smallest suboptimality gap $1/\Delta$, our bound reveals the relationship between the regret and the suboptimality gap of all arms. In addition, these works require prior knowledge of the independence number $\alpha$, which is NP-hard to compute, whereas SOG-BARBAT does not need such prior knowledge.
Remark 8.
Compared with lu2021stochastic, which adapts BARBAR to graph bandits with undirected graphs, our SOG-BARBAT method improves the corruption-dependent term from $O(\alpha C)$ to $O(C)$. Also, our method can be applied to bandits with directed graphs.